# Near-limit quantum control beyond analytic tractability in many-body spin systems

### v2

@@
Near-limit quantum control beyond analytic tractability in many-body spin systems
@@ Introduction
As quantum technologies transition from proof-of-principle demonstrations toward engineered platforms, control performance increasingly enters regimes in which previously negligible effects become consequential (?, ?, ?, ?). Assumptions that enable closed-form reasoning become progressively misaligned with weak residual effects that are omitted in tractable models yet dominate near-limit performance (?, ?). This structural tension motivates computational methods capable of relaxing assumptions required for analytic tractability, allowing them to function as soft guidance rather than hard design constraints.
@@
Computational approaches have enabled substantial progress across quantum technologies (?, ?), particularly for tasks with well-defined objectives and structured, learnable solution spaces (?, ?, ?, ?). However, certain central problems do not naturally meet these conditions and continue to resist systematic computational treatment, instead predominantly relying on closed-form analysis, notably pulse sequences that engineer time-averaged quantum dynamics (?, ?, ?). This resistance reflects an underlying circularity: assumptions introduced to ensure analytic tractability also define available search strategies, thereby preventing them from relaxing these very assumptions (?, ?, ?).
@@
Historically, it has been proven remarkably effective to focus attention on analytically tractable subspaces within high-dimensional design landscapes, capture dominant effects, and thus yield constructions approaching performance limits (?, ?, ?, ?). Outside these subspaces, high-performing solutions are often sparse and performance typically varies sharply under small design changes, rendering naive computational exploration ineffective. In near-limit regimes, however, the tractability-imposed boundaries become overly restrictive, by construction excluding unconventional solutions that may better capture weak, performance-limiting effects. Analytic tractability thus shifts from a guiding principle to a burden.
@@
Here, we address this challenge in pulse-sequence design for many-body spin control, a paradigmatic discrete and combinatorial optimization problem long constrained by analytic tractability. Since Hahn’s seminal work in 1950, pulse-sequence design has relied on average Hamiltonian theory (AHT) as its analytic foundation (?, ?), underpinning progress from NMR and MRI to a workhorse control technique across quantum platforms (?, ?, ?). Although its limitations have long been recognized (?, ?), pulse-sequence design has remained confined to the tractable subspaces defined by AHT-based structural constraints and screening rules (Fig. 1c). This reflects the same circularity: search strategies are formulated within theoretical frameworks, thereby reinforcing confinement to analytically tractable subspaces.
@@
To break this circularity, we repurpose assumptions required for analytic tractability from hard design constraints into soft guidance, and combine them with a stochastic tree search designed for general complex systems (?). Using this physics-guided computational framework (Fig. 1d), we identify a broad class of high-performing pulse sequences that consistently depart from long-standing analytic design principles (?, ?). Experiments in a solid-state spin ensemble confirm substantial coherence enhancements over analytically optimized baselines (Fig. 1a,e). Importantly, these unconventional sequences uncover interpretable structural features that can reliably predict performance, enabling rapid neural network–based evaluation. The efficiency gain allows expansion of the control alphabet from 8 symmetry-restricted pulses to 26,400 hardware-resolved options, despite the combinatorial scaling of the sequence space (Fig. 1f). Within such expanded spaces, fine-grained sequence variation provides the control resolution needed to resolve the impact of weak, hardware-specific effects in a controlled manner. Together, these results establish a qualitative shift beyond decades of traditional sequence-design approaches, and highlight a shared problem structure: near performance limits in realistic quantum hardware, control performance can be constrained by assumptions introduced to preserve analytic tractability.
Results
Analytic Performance Limits
@@ Results
We study a dense ensemble of approximately interacting spins in diamond (see Methods; Fig. 1a). Strong disorder, dipolar couplings, and control imperfections give rise to intrinsically complex many-body dynamics in the presence of environmental decoherence (Fig. 1b). The control objective is coherence preservation, quantified by the lifetime of prepared spin states under these conditions (see Methods).
@@
To establish the analytic performance frontier, we benchmark a hierarchy of analytically constructed baseline sequences composed of Clifford and pulses. Ramsey measurements yield an unprotected decay rate of approximately 500 kHz. The XY family suppresses this to about 40 kHz, while the analytically optimized DROID family further reduces it to roughly 10 kHz (?, ?, ?, ?). Continuous spin locking achieves decay rates approaching 1 kHz, setting a practical thermal floor (see Methods; Fig. 1e).
@@
To quantify the origins of this gap and establish a reference for subsequent optimization, we construct a coarse-grained decoherence budget and a calibrated simulator. The decoherence budget (Fig. 1e) assumes independent and additive decoherence contributions, following standard practice (?, ?, ?). It shows that approximately 98% of the total decay arises from disorder and dipolar interactions, identifying the dominant channels targeted by analytic design principles (see Methods; Fig. 1c). In parallel, the calibrated simulator reproduces baseline performance trends with high fidelity (Pearson correlation 0.93, Spearman correlation 0.95; Fig. 2a; see Methods).
Near-Limit Stringent Benchmark
The analytically optimized DROID family is designed to cancel all leading-order contributions from disorder, dipolar interactions, and pulse imperfections under the framework of AHT (?, ?, ?). These sequences do achieve strong experimental suppression of decoherence by mitigating the dominant decay channels. The residual 10 kHz decoherence is governed by effects that are either omitted or substantially simplified, and therefore lie beyond the descriptive capability of AHT and reduced models.
@@
Beyond Analytic Performance Limits
@@
Within DOESS, AHT-based criteria are repurposed as coarse, physically motivated filters, replacing the overly restrictive binary selection rules of traditional analytic design. Quantitative performance is then evaluated through simulation, free from the assumptions required for closed-form analysis (see Methods). To mitigate model–experiment mismatch and improve robustness, multiple optimization rounds are performed in parallel, each guided by a distinct and plausible simulator (Fig. 1d).
The integration of rapid rejection with robust ensemble evaluation enables our framework to efficiently navigate beyond analytically tractable subspaces defined by AHT (see Methods). We define a combinatorial search space of length-24 sequences composed of 13 discrete rotations (, , and about the or axes, as well as a no-pulse operation), and relax net-rotation constraints typically imposed for analytic convenience. These choices deliberately expand the search beyond symmetry-restricted constructions favored by analytic tractability. The inclusion of non-Clifford pulses generates substantially more complex spin trajectories than Clifford-only constructions (Fig. 1f).
@@
Experimentally, substantially larger relative improvements are observed, which is up to 150% (Fig. 2c), reflecting a pronounced mismatch between reduced models and experimental reality. Such experimental improvement arises from comparable signal contrast combined with a substantially extended coherence time (Fig. 2d,e). The corresponding raw data are shown in Fig. 2f, where the longest coherence times increase from approximately 100 to over 200 , or a reduction in the decay rate from 10 kHz to 5 kHz (see Methods).
@@
Hidden Structure Beyond Analytic Intuition
Beyond performance gains, our results reveal a structural mismatch between long-standing analytic intuition and actual near-limit control performance. Sequences deemed poorly optimized under standard single-cycle diagnostics can nevertheless perform exceptionally well under periodic repetition. This discrepancy indicates that the limitation lies not in the control protocols themselves, but in the diagnostic scale imposed by analytically convenient single-cycle reasoning. Specifically, DOESS-discovered sequences systematically deviate from the exact leading-order cancellation conditions enforced by AHT-based design (?, ?). These unconventional yet high-performing sequences illustrate how analytic tractability can shift from an effective guiding principle to a limiting constraint.
Under standard AHT-based intuition, the large single-cycle disorder residuals observed here would be expected to dominate decoherence (see Methods). According to the AHT-based decoherence budget, disorder dominates the decay rate (approximately 92%; Fig. 1e), and single-cycle AHT indicators heuristically suggest residual decay rates on the order of several tens of kilohertz (see Methods). Experimentally, however, the majority of these sequences exhibit decay rates around 10 kHz (Fig. 2f), substantially lower than this estimate.
Further analysis shows that this apparent mismatch arises because performance is governed by multi-cycle, rather than single-cycle dynamics. Relaxing net-rotation constraints enables non-identity sequences, for which disorder contributions are determined by the periodically repeated cycle rather than by the single-cycle evolution. For identity cycles, single-cycle and multi-cycle AHT analyses coincide. By contrast, for non-identity sequences accessed by DOESS, repetition defines effective cycles whose time-averaged disorder indicators converge upon repetition and approach values consistent with AHT predictions (see Methods).
@@
The physical relevance of multi-cycle analysis is further validated through predictive surrogate model training. When trained using single-cycle AHT indicators as input features, the surrogate achieves poor predictive accuracy (), effectively no better than random guessing. By contrast, when trained using indicator series that are computed across repeated cycles, the surrogate attains on randomly generated sequences (Fig. 3b). In addition, using indicator series as input features, t-SNE visualization (?) clearly separates computationally discovered sequences, analytic baselines, and random sequences (Fig. 3c). These physics-guided computational representations encode weak residual effects in a predictive and structured manner that departs from conventional closed-form descriptions.
Control Design at Hardware Resolution
The newly acquired computational capability enables control design at hardware resolution, where increased design freedom allows control protocols to better accommodate the full complexity of realistic hardware. At the hardware-resolved level, each control pulse is specified by its rotation axis and angle, both discretized in steps of , yielding over 26,000 distinct pulse operations (Fig. 4b; see Methods).
This combinatorial explosion was initially intractable, and the discovery stage was therefore restricted to shorter sequences with a reduced control alphabet consisting of Clifford operations augmented by four rotations. While access to pulse variations at hardware resolution introduces fine-grained control that mitigates single-pulse sensitivity and yields smoother performance landscapes under hardware-specific noise, the corresponding combinatorial growth renders direct application of DOESS at hardware resolution computationally infeasible (Fig. 4d).
@@ Discussion
Acknowledgements: The authors thank Haixin Qiu and Di Xu for valuable feedback that helped improve the manuscript. Y.Wei and B.P. acknowledge support from CityUHK start-up grant (No.9382006). This work is supported by the EU Horizon Project AMADEUS (101080136); SPINUS (101135699); EU Quantum Flagship Project C-QuENS (101135359); German Federal Ministry of Education and Research (BMBF) through DiaqNOS (13N16462), NeuroQ (13N16487), QSOLID (13N16159), QECHQS (16KIS1590K), and through the Future Cluster QSens QHMI2 (03ZU2110FB); and Deutsche Forschungsgemeinschaft (DFG) through projects GRK 2642, FOR 2724 (No. 384846402).
Author contributions: Y. Wang, Y. Wei, and J. Z. designed the project; R. P., J. Z., and Y. Wang conceived the initial idea; Y. Wei and J. W. managed the project; J. W. provided strategic guidance; Y. Wang developed the numerical simulator; Y. Wei and Y. Wang proposed the ensemble-simulation strategy; Y. Wei and B. P. conducted sequence optimizations and machine learning model training with contribution from H. P.; J. Z. and C. K. C. designed and built the experimental setup, conducted the experiments, and collected all data; J. Z. and Y. Wang analyzed the sequence structures; Y. Wang proposed the idea of physics-informed predictive modeling; B. P. generated the final figures with contributions from J. Z., and G. B.; Y. Wang, J. Z., and B. P. wrote the Supplementary Information; Y. Wang and Y. Wei prepared the initial draft with input from all authors; Y. Wang led the manuscript writing, including rewriting and integration of all components; Y. Wang and Y. Wei supervised the project.
Competing interests: Authors declare that they have no competing interests.
Code availability: Code and data are available at:
https://github.com/Bop2000/DOESS
Methods
Experimental Setup and Theoretical Modeling
The diamond sample contains a 1 -thick NV-rich layer with an concentration of approximately 1 ppm and an isotopic purity of 99.999% (Element Six Inc.). Optical spin initialization and fluorescence readout are performed using a confocal microscope with a focused laser spot of roughly 1 in diameter. The corresponding detection volume of 3 contains approximately NV centers. Such many-body spin systems serve as paradigmatic platforms for quantum sensing (?, ?, ?, ?) and simulation (?, ?, ?, ?).
@@ Methods
Spin Coherence and Score
Spin coherence is quantified as the arithmetic average of the survival probabilities of spin states initially polarized along the , , and axes, represented by the state vectors , , and , respectively (?). This definition provides a unified metric applicable across different control protocols, including Ramsey and spin-locking measurements.
@@
To obtain compact performance descriptors, the coherence functions from Eq. (1) are fit to a single-exponential decay model, . Here, is the contrast and is the decay rate along the axis . Averaging over the three orthogonal axes yields an overall decay rate , and the coherence time is defined as , following Ref. (?).
@@
Decoherence Budget and Calibrated Simulator
We consider a set of analytically constructed baseline sequences, including XY8/XY16, DROID, and continuous spin locking, to characterize dominant decoherence mechanisms and to calibrate a numerical simulator. XY8/XY16 sequences primarily suppress disorder-induced dephasing, while DROID further mitigates contributions from dipolar spin-spin interactions and exhibits strong robustness to pulse imperfections. In continuous spin-locking experiments, spins are initialized along an axis in the plane and driven continuously by a resonant microwave field. Under the secular approximation (?), perturbations that do not commute with the driving field—including disorder, dipolar interactions, and environmental noise—are strongly suppressed, defining a practical coherence limit. Detailed descriptions are provided in the Supplementary Information: Experimental Details.

### v3

@@
Abstract
@@ I Introduction
As quantum technologies transition from proof-of-principle demonstrations toward engineered platforms, control performance increasingly enters regimes in which previously negligible effects become consequential [21, 4, 1, 8]. Assumptions that enable closed-form reasoning become progressively misaligned with weak residual effects that are omitted in tractable models yet dominate near-limit performance [29, 9]. This structural tension motivates computational methods capable of relaxing assumptions required for analytic tractability, allowing them to function as soft guidance rather than hard design constraints.
@@
Computational approaches have enabled substantial progress across quantum technologies [9, 31], particularly for tasks with well-defined objectives and structured, learnable solution spaces [7, 39, 15, 18]. However, certain central problems do not naturally meet these conditions and continue to resist systematic computational treatment, instead predominantly relying on closed-form analysis, notably periodic pulse sequences that engineer time-averaged quantum dynamics [14, 13, 27]. This resistance reflects an underlying circularity: assumptions introduced to ensure analytic tractability also define available search strategies, thereby preventing them from relaxing these very assumptions [32, 9, 26].
@@
Historically, it has been proven remarkably effective to focus attention on analytically tractable subspaces within high-dimensional design landscapes, capture dominant effects, and thus yield constructions approaching performance limits [20, 34, 2, 33]. Outside these subspaces, high-performing solutions are sparse, and performance is highly sensitive to small design changes, which often causes conventional gradient-based algorithms to become trapped at suboptimal points [30, 3]. In near-limit regimes, however, the tractability-imposed boundaries become overly restrictive, by construction excluding unconventional solutions that may better capture weak, performance-limiting effects. Analytic tractability thus shifts from a guiding principle to a burden.
@@
Here, we address this challenge in pulse-sequence design for many-body spin control, a paradigmatic discrete and combinatorial optimization problem long constrained by analytic tractability. Since Hahn’s seminal work in 1950, pulse-sequence design has relied on average Hamiltonian theory (AHT) as its analytic foundation [20, 34], underpinning progress from NMR and MRI to a workhorse control technique across quantum platforms [13, 28, 11]. Although its limitations have long been recognized [19, 25], pulse-sequence design has remained confined to the tractable subspaces defined by AHT-based structural constraints and screening rules (Fig. 1c). This reflects the same circularity: search strategies are formulated within theoretical frameworks, thereby reinforcing confinement to analytically tractable subspaces.
@@
To break this circularity, we repurpose assumptions required for analytic tractability from hard design constraints into soft guidance, and combine them with a stochastic tree search designed for general complex systems [36]. Using this physics-guided computational framework (Fig. 1d), we identify a broad class of high-performing pulse sequences that consistently depart from long-standing analytic design principles [34, 13]. Experiments in a solid-state spin ensemble confirm substantial coherence enhancements over analytically optimized baselines (Fig. 1a,e). Importantly, these unconventional sequences uncover interpretable structural features that can reliably predict performance, enabling rapid neural network–based evaluation. The efficiency gain allows expansion of the control alphabet from 8 symmetry-restricted pulses to 26,400 hardware-resolved options, despite the combinatorial scaling of the sequence space (Fig. 1f). Within such expanded spaces, fine-grained sequence variation provides the control resolution needed to resolve the impact of weak, hardware-specific effects in a controlled manner.
These results establish a qualitative shift beyond decades of traditional sequence-design approaches, and highlight a shared problem structure: near performance limits in realistic quantum hardware, control performance can be constrained by assumptions introduced to preserve analytic tractability.
II Analytic Performance Limits
@@
We study a dense ensemble of approximately interacting spins in diamond (see Appendix; Fig. 1a). Strong disorder, dipolar couplings, and control imperfections give rise to intrinsically complex many-body dynamics in the presence of environmental decoherence (Fig. 1b). The control objective is coherence preservation, quantified by the lifetime of prepared spin states under these conditions (see Appendix).
@@
To establish the analytic performance frontier, we benchmark a hierarchy of analytically constructed baseline sequences composed of discrete and pulses. Ramsey measurements yield an unprotected decay rate of approximately 500 kHz. The XY family suppresses this to about 40 kHz, while the analytically optimized DROID family further reduces it to roughly 10 kHz [34, 13, 17, 38]. Continuous spin locking achieves decay rates approaching 1 kHz, setting a practical thermal floor (see Appendix; Fig. 1e).
@@
To quantify the origins of this gap and establish a reference for subsequent optimization, we construct a coarse-grained decoherence budget and a calibrated simulator. The decoherence budget (Fig. 1e) assumes independent and additive decoherence contributions, following standard practice [5, 22, 37]. It shows that approximately 98% of the total decay arises from disorder and dipolar interactions, identifying the dominant channels targeted by analytic design principles (see Appendix; Fig. 1c). In parallel, the calibrated simulator reproduces baseline performance trends with high fidelity (Pearson correlation 0.93, Spearman correlation 0.95; Fig. 2a; see Appendix).
III Near-Limit Stringent Benchmark
The analytically optimized DROID family is designed to cancel all leading-order contributions from disorder, dipolar interactions, and pulse imperfections under the framework of AHT [13, 17, 38]. These sequences do achieve strong experimental suppression of decoherence by mitigating the dominant decay channels. The residual 10 kHz decoherence is governed by effects that are either omitted or substantially simplified, and therefore lie beyond the descriptive capability of AHT and reduced models.
@@ III Near-Limit Stringent Benchmark
IV Beyond Analytic Performance Limits
@@
Within DOESS, AHT-based criteria are repurposed as coarse, physically motivated filters, replacing the overly restrictive binary selection rules of traditional analytic design. Quantitative performance is then evaluated through simulation, free from the assumptions required for closed-form analysis (see Appendix). To mitigate model–experiment mismatch and improve robustness, multiple optimization rounds are performed in parallel, each guided by a distinct and plausible simulator (Fig. 1d).
The integration of rapid rejection with robust ensemble evaluation enables our framework to efficiently navigate beyond analytically tractable subspaces defined by AHT (see Appendix). We define a combinatorial search space of length-24 sequences composed of 13 discrete rotations (, , and about the or axes, as well as a no-pulse operation), and relax net-rotation constraints typically imposed for analytic convenience. These choices deliberately expand the search beyond symmetry-restricted constructions favored by analytic tractability. The inclusion of non-Clifford pulses generates substantially more complex spin trajectories than Clifford-only constructions (Fig. 1f).
@@
Experimentally, substantially larger relative improvements are observed, which is up to 150% (Fig. 2c), reflecting a pronounced mismatch between reduced models and experimental reality. Such experimental improvement arises from comparable signal contrast combined with a substantially extended coherence time (Fig. 2d,e). The corresponding raw data are shown in Fig. 2f, where the longest coherence times increase from approximately 100 to over 200 , or a reduction in the decay rate from 10 kHz to 5 kHz (see Appendix).
@@
V Hidden Structure Beyond Analytic Intuition
Beyond performance gains, our results reveal a structural mismatch between long-standing analytic intuition and actual near-limit control performance. Sequences deemed poorly optimized under standard single-cycle diagnostics can nevertheless perform exceptionally well under periodic repetition. This discrepancy indicates that the limitation lies not in the control protocols themselves, but in the diagnostic scale imposed by analytically convenient single-cycle reasoning. Specifically, DOESS-discovered sequences systematically deviate from the exact leading-order cancellation conditions enforced by AHT-based design [34, 13]. These unconventional yet high-performing sequences illustrate how analytic tractability can shift from an effective guiding principle to a limiting constraint.
Under standard AHT-based intuition, the large single-cycle disorder residuals observed here would be expected to dominate decoherence (see Appendix). According to the AHT-based decoherence budget, disorder dominates the decay rate (approximately 92%; Fig. 1e), and single-cycle AHT indicators heuristically suggest residual decay rates on the order of several tens of kilohertz (see Appendix). Experimentally, however, the majority of these sequences exhibit decay rates around 10 kHz (Fig. 2f), substantially lower than this estimate.
Further analysis shows that this apparent mismatch arises because performance is governed by multi-cycle, rather than single-cycle dynamics. Relaxing net-rotation constraints enables non-identity sequences, for which disorder contributions are determined by the periodically repeated cycle rather than by the single-cycle evolution. For identity cycles, single-cycle and multi-cycle AHT analyses coincide. By contrast, for non-identity sequences accessed by DOESS, repetition defines effective cycles whose time-averaged disorder indicators converge upon repetition and approach values consistent with AHT predictions (see Appendix).
@@ V Hidden Structure Beyond Analytic Intuition
The physical relevance of multi-cycle analysis is further validated through predictive surrogate model training. When trained using single-cycle AHT indicators as input features, the surrogate achieves poor predictive accuracy (), effectively no better than random guessing. By contrast, when trained using indicator series that are computed across repeated cycles, the surrogate attains on randomly generated sequences (Fig. 3b). In addition, using indicator series as input features, t-SNE visualization [24] clearly separates computationally discovered sequences, analytic baselines, and random sequences (Fig. 3c). These physics-guided computational representations encode weak residual effects in a predictive and structured manner that departs from conventional closed-form descriptions.
VI Control Design at Hardware Resolution
The newly acquired computational capability enables control design at hardware resolution, where increased design freedom allows control protocols to better accommodate the full complexity of realistic hardware. At the hardware-resolved level, each control pulse is specified by its rotation axis and angle, both discretized in steps of , yielding over 26,000 distinct pulse operations (Fig. 4b; see Appendix).
This combinatorial explosion was initially intractable, and the discovery stage was therefore restricted to shorter sequences with a reduced control alphabet consisting of Clifford operations augmented by four rotations. While access to pulse variations at hardware resolution introduces fine-grained control that mitigates single-pulse sensitivity and yields smoother performance landscapes under hardware-specific noise, the fundamental challenge of a highly rugged design space persists. The corresponding combinatorial growth renders direct application of DOESS at hardware resolution computationally infeasible (Fig. 4d). More critically, this high-dimensional space remains non-convex and densely populated with local optimum traps. Consequently, gradient-based optimization algorithms are prone to becoming trapped in these suboptimal points, failing to navigate the complex performance topography towards the global optimum.
@@ VII Discussion
Acknowledgments
The authors thank Haixin Qiu and Di Xu for valuable feedback that helped improve the manuscript. Y.Wei and B.P. acknowledge support from CityUHK start-up grant (No.9382006). This work is supported by the EU Horizon Project AMADEUS (101080136); SPINUS (101135699); EU Quantum Flagship Project C-QuENS (101135359); German Federal Ministry of Education and Research (BMBF) through DiaqNOS (13N16462), NeuroQ (13N16487), QSOLID (13N16159), QECHQS (16KIS1590K), and through the Future Cluster QSens QHMI2 (03ZU2110FB); and Deutsche Forschungsgemeinschaft (DFG) through projects GRK 2642, FOR 2724 (No. 384846402).
Y. Wang, Y. Wei, and J. Z. designed the project; R. P., J. Z., and Y. Wang conceived the initial idea; Y. Wei and J. W. managed the project; J. W. provided strategic guidance; Y. Wang developed the numerical simulator; Y. Wei and Y. Wang proposed the ensemble-simulation strategy; Y. Wei and B. P. conducted sequence optimizations and machine learning model training with contribution from H. P.; J. Z. and C. K. C. designed and built the experimental setup, conducted the experiments, and collected all data; J. Z. and Y. Wang analyzed the sequence structures; Y. Wang proposed the idea of physics-informed predictive modeling; B. P. generated the final figures with contributions from J. Z., and G. B.; Y. Wang, J. Z., and B. P. wrote the Supplementary Information; Y. Wang and Y. Wei prepared the initial draft with input from all authors; Y. Wang led the manuscript writing, including rewriting and integration of all components; Y. Wang and Y. Wei supervised the project.
Data and code availability
All data and code are available as a GitHub repository: https://github.com/Bop2000/DOESS
Appendix A Experimental Setup and Theoretical Modeling
The diamond sample contains a 1 -thick NV-rich layer with an concentration of approximately 1 ppm and an isotopic purity of 99.999% (Element Six Inc.). Optical spin initialization and fluorescence readout are performed using a confocal microscope with a focused laser spot of roughly 1 in diameter. The corresponding detection volume of 3 contains approximately NV centers. Such many-body spin systems serve as paradigmatic platforms for quantum sensing [6, 5, 10, 16] and simulation [12, 40, 22, 23].
@@ Appendix A Experimental Setup and Theoretical Modeling
Appendix B Spin Coherence and Score
Spin coherence is quantified as the arithmetic average of the survival probabilities of spin states initially polarized along the , , and axes, represented by the state vectors , , and , respectively [13]. This definition provides a unified metric applicable across different control protocols, including Ramsey and spin-locking measurements.
@@ Appendix B Spin Coherence and Score
To obtain compact performance descriptors, the coherence functions from Eq. (1) are fit to a single-exponential decay model, . Here, is the contrast and is the decay rate along the axis . Averaging over the three orthogonal axes yields an overall decay rate , and the coherence time is defined as , following Ref. [13].
@@
Appendix C Decoherence Budget and Calibrated Simulator
We consider a set of analytically constructed baseline sequences, including XY8/XY16, DROID, and continuous spin locking, to characterize dominant decoherence mechanisms and to calibrate a numerical simulator. XY8/XY16 sequences primarily suppress disorder-induced dephasing, while DROID further mitigates contributions from dipolar spin-spin interactions and exhibits strong robustness to pulse imperfections. In continuous spin-locking experiments, spins are initialized along an axis in the plane and driven continuously by a resonant microwave field. Under the secular approximation [35], perturbations that do not commute with the driving field—including disorder, dipolar interactions, and environmental noise—are strongly suppressed, defining a practical coherence limit. Detailed descriptions are provided in the Supplementary Information: Experimental Details.